OpenStudy (anonymous):

1.you pay 1.75 for 5 bagels. what is the unit price? a. 0.09 b. 0.45 c. 0.35 d. 0.44 2. which is the best buy? a. 12.60 for 7 bottles of juice b. 10.98 for 6 bottles of juice c. 18.10 for 10 bottles of juice d. 16.38 for 9 bottles of juice

OpenStudy (anonymous):

plz help me !!!!!!!!!!!

sammixboo (sammixboo):

#1 just take 1.75 and divide it by 5, so what is 1.75/5? #2 a - 12.60/7 b - 10.98/6 c - 18.10/10 d - 16.38/9 Which ever one has the lowest value is yur answer

OpenStudy (anonymous):

For number one, it is simply just taking the price the customer payed, and dividing it by the amount of bagels purchased, which would be 1.75 divided by 5, 1.75/5 = 0.35. Which would make "C" your answer for number one.

OpenStudy (anonymous):

And for number two, the best deal would be the price with the lowest cost per unit. So 12.60/7 = 1.8$ 10.98/6 = 1.83$ 18.10/10 = 1.81$ 16.38/9 = 1.82$ Therefore, the answer for number two would be "A"